Determine the x-coordinate of the centroid of the shaded area. - - - - - - 1 1
Answer #1

4 soon! itd y2 2² =1 92 42-4, y? - 1-H? 1 (77,5). a 2 ५। dol --> 90 3 * / ** 0 Y = b-kost a = dA b a da = (d2-8,) aset dĄ = [A 1 Ha,Lde = Jae-sado n=0년 Q2 42 let a ² n² = 4 _2nde = de 2 ) non=-2 / - S7 발 = (if - 9 3 to → + = Q2. M9 ale - [ ] x = 3